# Scaling Laws for Pre-training Agents and World Models

### v1

@@ 4.1 Scaling analysis in world modeling
Figures 3 & 4 present our results for the task of world modeling, with the scaling law coefficients summarised in Table 1. For WM-Token-256 we find that the optimal coefficients for model and dataset size are both $\approx 0.5$, e.g. one should increase both model and dataset size in the same proportions. This matches the scaling laws observed in LLMs [Hoffmann et al., 2022]. Increasing the number of tokens per image to $540$ for WM-Token-540 changes the optimal trade-off between model and dataset size, skewing towards model size; $N_{\text{optimal}}=0.62$, $D_{\text{optimal}}=0.37$. We discuss this further in Section 5.3.
@@ 4.2 Scaling analysis in behavior cloning
To better understand the change in the scaling law coefficients, we now consider the BC-CNN architecture for the task of BC in Figure 6. For this architecture, we observe that the coefficients now skew towards model size (similarly to those in [Tuyls et al., 2023]), with $N_{\text{optimal}}=0.66$, and $D_{\text{optimal}}=0.34$. Section 5.2 provides more intuition on the differences between the WM-Token and BC-Token setups that lead to this change.
@@ 5 Further analysis
  • Q3: Why does increasing the amount of tokens per image observation (256 to 540) lead to an increase in the optimal model size coefficient (0.49 to 0.62)? (Section 5.3)
@@
$N_{\text{optimal}}\propto C^{0.63}$                      $N_{\text{optimal}}\propto C^{0.50}$                      $N_{\text{optimal}}\propto C^{0.15}$
@@
For further insight into how compression affects the optimal model size coefficient, we ran a small scale experiment in language modeling using two text representations; 1) ASCII character-level tokenization. (low compression) 2) GPT-2 tokenizer (high compression). We used the BookCorpus dataset [Zhu et al., 2015], and trained models past their compute-optimal point, so the Frontier fit method could be used for coefficient estimation.
Appendix C shows results. Under the character-level tokenizer (low compression), we find $N_{\text{optimal}}\propto C^{0.66}$. For the GPT-2 tokenizer (high compression), we find $N_{\text{optimal}}\propto C^{0.44}$. The key takeaway is that the more compressed representation has led to a lower optimal model size coefficient, which reflects what we see in the world modeling experiments.
@@ 6 Discussion & conclusion
Ethics Statement
Data for this project was provided via a partnership with Ninja Theory, who collected a large corpus of human gameplay data for their game Bleeding Edge. Data collection was covered by an End User License Agreement (EULA) and our use of the data was governed by a data sharing agreement with the game studio, and approved by our institution’s IRB. This data was recorded between September 2020 and October 2022. To minimize risk to human subjects, any personally identifiable information (Xbox user ID) was removed from the data. The resulting data was cleaned to remove errors and data from bad actors.
Appendix A Extended related work
Here we contrast more granular details of several related works.
Tuyls et al. [2023] provide (amongst other things) a valuable initial study on the relationship between compute and pre-training loss for BC. They focus on Atari and Nethack games, where the dataset is generated by a fixed pre-trained agent. They test one architecture – a single-layer LSTM network, scaling width-wise to increase parameter count. They find the model size coefficient $N_{\text{optimal}}\propto C^{a}$ varies between 0.58 to 0.79 for different games.
By contrast, our work focuses on transformer-based models, and we consider two variants of architecture. We also train on human behavior, rather than pre-trained agents (which could be viewed as model distillation). Additionally, we go beyond BC and also study scale in world modeling.
Scaling laws in other domains. Scaling laws have also been observed in auto-regressive modeling of other modalities. Most relevant to our world modeling task are those of video and images. Henighan et al. [2020] found the optimal trade off between model and dataset size to match their reported LLM coefficient ($N_{\text{optimal}}\propto C^{0.7}$) and was not affected by tokenizer. Our experiments offer different findings in the domain of world modeling on human data – we use updated methodologies to measure this trade-off [Hoffmann et al., 2022], finding it is affected by the tokenizer. Tian et al. [2024] show scaling laws emerge in image modeling with non raster-ordered next-token prediction.
Appendix B Scaling experiments further details
B.1 Hyperparameters
We trained three VQGANs from scratch with reconstruction losses.
  • BE-Small. Based on Esser et al. [2021], uses $d_{z}=256$, $V_{o}=4096$, $h=w=128$, with 28M parameters, and a CNN design. It was trained on a single SkyGarden Bleeding Edge map.
  • BE-Large. Based on Yu et al. [2022], uses $d_{z}=540$, $V_{o}=4096$, $h=180$, $w=300$, with 150M parameters, and a vision transformer design. It was trained on all seven Bleeding Edge maps.
BC-CNN details. We use $h=w=128$. The $0.6M$ paramter CNN is similar to that used by [Baker et al., 2022], however it uses ConvNext blocks [Liu et al., 2022]. The CNN produces an embedding of size $1024$ which is then put through a linear layer to obtain a vector matching the transformer’s embedding dimension.
Transformer configurations are given in Table 2. We describe the parameters for the WM-Token architecture. Note that MLP layers are four times the width of embed dim.
@@ B.1 Hyperparameters
B.2 Training details
All transformers are trained with a variant of nanoGPT [Karpathy, 2022] using PyTorch Lightning [Falcon and The PyTorch Lightning team, 2019].
@@ B.3 Dataset details
B.3.1 On the infinite data regime
@@
WM-Token-540, BC-Token-540. We trained on the 7 maps dataset, with 1.63B observation-action pairs. Models used the tokenized architecture with the large VQGAN, so each observation-action pair creates $540+16=556$ transformer inputs, for a total of 1.63B×556=906$\times$ 556=906× 556 = 906B training tokens. Muennighoff et al. [2024] observe that tokens may be reused up to four times with negligible departure from the infinite data regime. This produces 3.6T tokens. For a 200M parameter model the compute allowed by the infinite data regime is $C=6ND=6\times 200\text{M}\times 3.6\text{T}=4.3\times 10^{21}$ FLOPs.
@@
Appendix C Further analysis details
@@
Figure 10: Relating to Section 5.3, character-level (low compression). Utilising the frontier fit (middle and right) we derive the power law coefficient for $N_{\text{optimal}}$ as $0.66$.
Figure 11: Relating to Section 5.3, GPT-2 tokenizer (high compression). Utilising the frontier fit (middle and right) we derive the power law coefficient for $N_{\text{optimal}}$ as $0.44$, an increase from $0.66$ in Figure 10 found when utilising a lower compression character-level tokenizer.

### v2

@@ 4.1 Scaling analysis in world modeling
Figures 6 & 7 present our results for the task of world modeling, with the scaling law coefficients summarized in Table 1. For WM-Token-256 we find that the optimal coefficients for model and dataset size are both $\approx 0.5$, e.g. one should increase both model and dataset size in the same proportions. This matches the scaling laws observed in LLMs (hoffmann2022training). Increasing the number of tokens per image to $540$ for WM-Token-540 changes the optimal trade-off between model and dataset size, skewing towards model size; $N_{\text{optimal}}=0.62$, $D_{\text{optimal}}=0.37$. We discuss this further in Section 5.3.
Appendix Figure 16 visualizes the power law fits for the RT-1 robotics dataset, confirming that this predictable scaling behavior is not specific to human behavior in video games, and also emerges on real-world robotics tasks with high-skill human operators.
@@ 4.2 Scaling analysis in behavior cloning
To better understand the change in the scaling law coefficients, we now consider the BC-CNN architecture for the task of BC in Figure 9. For this architecture, we observe that the coefficients now skew towards model size (similarly to those in (tuyls2023scalingimitation)), with $N_{\text{optimal}}=0.66$, and $D_{\text{optimal}}=0.34$. Section 5.2 provides more intuition on the differences between the WM-Token and BC-Token setups that lead to this change.
@@ 5 Further analysis
  • Q3: Why does increasing the tokens per image observation (256 to 540) lead to an increase in the optimal model size coefficient (0.49 to 0.62)? (Section 5.3)
@@
[…]
Refer to caption Refer to caption Refer to caption
@@
Using the small scale RT-1 dataset, we conduct a more thorough investigation of the effect of tokens-per-image observation on scaling coefficients. First we train a range of image tokenizers with $z_{o}\in[16,36,64,100,256]$, visualized in Figure 15. For each VQVAE, we then train a range of WM-Token model sizes $N\in[0.08M,0.2M,0.28M,0.54M,0.99M]$, and measure scaling coefficients using the frontier fit method, repeating three times.
Figure 12 plots all coefficient vs. tokens-per-image – we observe that the optimal parameter scaling coefficient increases with decreasing compression.
Figure 12: RT-1 experiments. Optimal parameter coefficient vs. number of tokens per observation, with three repeated runs per VQVAE.
To investigate whether compression affects the optimal model size coefficient outside of embodied domains, we ran a small scale experiment in language modeling using two text representations; 1) ASCII character-level tokenization. (low compression) 2) GPT-2 tokenizer (high compression). We used the BookCorpus dataset (Zhu_2015_ICCV), and trained models past their compute-optimal point, so the Frontier fit method could be used for coefficient estimation.
Appendix B shows results. Under the character-level tokenizer (low compression), we find $N_{\text{optimal}}\propto C^{0.66}$. For the GPT-2 tokenizer (high compression), we find $N_{\text{optimal}}\propto C^{0.44}$. Hence, in language the more compressed representation also leads to a lower optimal model size coefficient.
@@ 6 Discussion & conclusion
Ethics Statement. Data for this project was provided via a partnership with Ninja Theory, who collected a large corpus of human gameplay data for their game Bleeding Edge. Data collection was covered by an End User License Agreement (EULA) and our use of the data was governed by a data sharing agreement with the game studio, and approved by our institution’s IRB. This data was recorded between September 2020 and October 2022. To minimize risk to human subjects, any personally identifiable information (Xbox user ID) was removed from the data. The resulting data was cleaned to remove errors and data from bad actors.
[…]
The appendix is organized as follows.
  • Appendix A contains details on the training of the model configurations, hyperparameters, and a description of the datasets used.
  • Appendix B contains results from Section 5.3.
  • Appendix C contains further details on training world models on robotics.
  • Appendix D contains further results demonstrating the link between pre-training loss and performance.
Appendix A Scaling experiments further details
This section provides experimental details for all experiments on the primary Bleeding Edge dataset.
A.1 Hyperparameters
We trained two VQGANs from scratch with reconstruction losses.
  • BE-Small. Based on esser2021vqgan, uses $d_{z}=256$, $V_{o}=4096$, $h=w=128$, with 28M parameters, and a CNN design. It was trained on a single SkyGarden Bleeding Edge map.
  • BE-Large. Based on yu2022vqganvit, uses $d_{z}=540$, $V_{o}=4096$, $h=180$, $w=300$, with 150M parameters, and a vision transformer design. It was trained on all seven Bleeding Edge maps.
We selected the numbers of tokens per image based on qualitative assessment of reconstructions. We found that 256 tokens per image was the minimum that still allowed a reconstruction to capture the majority of salient gameplay details. However certain details still were lacking, such as an enemy player’s health bars – hence we also considered a 540 token version that provided a higher quality reconstruction.
BC-CNN details. We use $h=w=128$. The $0.6M$ paramter CNN is similar to that used by (baker2022vpt), however it uses ConvNext blocks (liu_convnet_2022). The CNN produces an embedding of size $1024$ which is then put through a linear layer to obtain a vector matching the transformer’s embedding dimension.
Transformer configurations are given in Table 2. We describe the parameters for the WM-Token architecture. Note that MLP layers are four times the width of embed dim. Model configurations roughly followed the model configurations used in Table A9 of hoffmann2022training, where residual stream dimension, number of layers, and number of heads were roughly increased proportionally.
@@ A.1 Hyperparameters
A.2 Training details
All transformers are trained with a variant of nanoGPT (Karpathy2022) using PyTorch Lightning (Falcon_PyTorch_Lightning_2019).
This section lists key hyperparameters. Note that it was important to find optimization settings that produced the lowest possible loss for a given model size. In general larger models require smaller learning rates. Our approach first optimized the smallest model through a grid sweep, we would then sequentially run a sweep over the next largest model, starting at the smaller model’s optimized learning rate. Table 3-6 provide final settings.
@@ A.3 Dataset details
A.3.1 Infinite data regime allowed FLOPs
@@
WM-Token-540, BC-Token-540. We trained on the 7 maps dataset, with 1.63B observation-action pairs. Models used the tokenized architecture with the large VQGAN, so each observation-action pair creates $540+16=556$ transformer inputs, for a total of 1.63B×556=906$\times$ 556=906× 556 = 906B training tokens. muennighoff2024repeatscaling observe that tokens may be reused up to four times with negligible departure from the infinite data regime. This produces 3.6T tokens. For a 200M parameter model the compute allowed by the infinite data regime is $C=6ND=6\times 200\text{M}\times 3.6\text{T}=4.3\times 10^{21}$ FLOPs.
@@
Appendix B Further analysis details
@@
Figure 13: Relating to Section 5.3, character-level (low compression). Utilising the frontier fit (middle and right) we derive the power law coefficient for $N_{\text{optimal}}$ as $0.66$ and $D_{\text{optimal}}$ as $0.34$.
Figure 14: Relating to Section 5.3, GPT-2 tokenizer (high compression). Utilising the frontier fit (middle and right) we derive the power law coefficient for $N_{\text{optimal}}$ as $0.44$ and $D_{\text{optimal}}$ as $0.56$, an increase from $0.66$ in Figure 13 found when utilising a lower compression character-level tokenizer.
Appendix C World modeling for robotics experimental details
This section provides experimental details for WM experiments on the secondary RT-1 dataset.
C.1 Dataset
We resized the RT-1 dataset to 128x128 pixels per image. For action labels, we take the 3D world_vector coordinates, combined with the 1D gripper_closedness_action vector, to make an action vector with four dimensions. All are in the range -1 to 1, and these are discretized into 500 evenly spaced buckets.
C.2 VQVAEs
We trained a set of five VQVAEs using the implementation from https://github.com/nadavbh12/VQ-VAE. We set $z_{o}\in[16,36,64,100,256]$ and $V_{o}=4096$, training each VQVAE for 40,000 updates on batches of 128. Reconstructions are visualized in Figure 15.
Figure 15: VQVAE reconstructions on the RT-1 dataset for differing numbers of tokens per observation, $z_{o}\in[16,36,64,100,256]$.
C.3 Transformer training details
Table 7 provides training details for the model sizes tested. Figure 16 shows one example set of training curves per VQVAE.
Table 7: Hyperparameters for WM-Token in RT-1 experiments.
$N$ Seq len Context length Tokens per update Learning rate
0.08M 2 2$(z_{o}+4)$ 34,000 0.01
0.2M 2 2$(z_{o}+4)$ 34,000 0.005
0.28M 2 2$(z_{o}+4)$ 34,000 0.004
0.54M 2 2$(z_{o}+4)$ 34,000 0.0027
0.99M 2 2$(z_{o}+4)$ 34,000 0.002
$z_{o}=16$, $N_{\text{optimal}}\propto C^{0.56}$, $N_{\text{optimal}}\propto D^{0.44}$
$z_{o}=36$, $N_{\text{optimal}}\propto C^{0.60}$, $N_{\text{optimal}}\propto D^{0.40}$
$z_{o}=64$, $N_{\text{optimal}}\propto C^{0.61}$, $N_{\text{optimal}}\propto D^{0.39}$
$z_{o}=100$, $N_{\text{optimal}}\propto C^{0.60}$, $N_{\text{optimal}}\propto D^{0.40}$
$z_{o}=256$, $N_{\text{optimal}}\propto C^{0.65}$, $N_{\text{optimal}}\propto D^{0.34}$
Figure 16: RT-1 experiments. Note that the optimal parameter coefficient increases with the number of tokens per observation.
Appendix D Pre-training loss vs. world modeling metrics
This section presents evidence for pre-training loss correlating with WM performance. We use metrics commonly used to assess the quality of the world models (yang2023unisim), originally developed in the video generation literature. Conditioned on an initial real frame and a sequence of real actions, we compare the observations generated by a world model, with the real sequence of observations, measuring FVD and LPIPS. Specifically, we generate 1024 videos each of 10 seconds. We perform this for various checkpoints on each size in our WM-Token-256 set of models. This allows a plot of the checkpoint pre-training loss vs video generation metric to be assessed.
Figure 3 shows results. We find correlations of 0.77, 0.83 for LPIPS and FVD respectively. Two early checkpoints from the 894M model are the only significant anomalies to trend of metrics improving with loss. This evidences the strong relationship between pre-training loss and world model quality.